OpenStudy (studytheworld):

Help please! Will medal! helppp! Suppose a triangle has two sides of lengths 33 and 37, and that the angle between these two sides is 120 degrees. What is the length of the third side of the triangle?

OpenStudy (studytheworld):

@TrojanPoem Help with this one?

OpenStudy (ihelpyoulearn):

@math&ing001 Do you think you could help?

OpenStudy (trojanpoem):

 use the cosine law: a^2 = b^2 + c^2 - 2 bc cosA a = sqrt( (33)^2 + (37)^2 - 2 * 33 * 37 * cos 120) = 60.65

OpenStudy (studytheworld):

Yay thank you so much! you help a lot

OpenStudy (trojanpoem):

Any time.